# Semi-analytical covariance matrices for two-point correlation function for DESI 2024 data

### v2

@@ 5.2 Shot-noise rescaling values
LRG pre-recon | 0.86 | 0.95 | ||
LRG post-recon | 0.85 | 0.96 | ||
ELG pre-recon | 0.68 | 0.72 | ||
ELG post-recon | 0.69 | 0.74 | ||
BGS pre-recon | 0.85 | 0.88 | ||
BGS post-recon | 0.87 | 0.92 |
North | South | |
LRG pre-recon | 0.97 | 0.99 |
LRG post-recon | 0.98 | 1.00 |
BGS pre-recon | 1.03 | 0.96 |
BGS post-recon | 0.96 | 1.02 |
As we pointed out in Section 2.1, our method uses a rescaling of the shot noise contribution to account for differences between the true small-scale contributions and our Gaussian approximation. After reaching a relatively uniform convergence level, we inspect the shot-noise rescaling values obtained from fitting the jackknife covariances in Table 2. Interestingly, we find that all the shot-noise rescaling values are smaller than 1, i.e., that the jackknife variations are smaller than predicted by Gaussian approximation with standard Poisson shot-noise (Eq. 2.4). This is different from what has been found in previous mock studies [26, 27, 28, 31], but we believe that this is a consequence of the effects of fiber assignment incompleteness. In contrast, the results for the early DESI data131313Early DESI data is not the Early Data Release [4], which had high completeness, but the data collected during the first two months of the main survey, thus included in DR1 [5]. [30], which had more uniform completeness, showed shot-noise rescaling values close to 1, as shown in Table 3.
Next, we see that the shot-noise rescaling values are significantly lower on mocks compared to the data. Again, the difference is most pronounced for ELG. This is an indication that the fast fiber assignment does not exactly reproduce the covariance matrix/shot-noise properties of the real one.
@@
In light of the concerns about fiber assignment and jackknife discussed in Section 4, these results warrant further investigation. We have come up with a self-consistency test for our covariance matrix models: optimization of the shot-noise rescaling based on mock sample covariance (detailed in Appendix C and particularly Fig. 4). We show the resulting values along with the baseline, jackknife-based ones in Table 10, and find them very close for all cases. In other words, calibration of shot-noise rescaling on jackknives still brings us close to an optimal fit on the mocks. With that, we have decided to proceed further with the validation process. This will show how close this nearly optimal fit is to the mock sample covariance.
5.3 Observable space
@@
We now begin to compare the RascalC results with the mock sample covariance matrices. We start in Table 4 with the widest available range: using all three multipoles. We remind that for a fixed tracer pre- or post-recon, all 10 single-mock results are compared to the same sample covariance matrix, so that noise, which is accounted for in the “perfect case” is “fixed” independently from mock-to-mock scatter in RascalC results. Ideally, we would like the comparison measures to be within the perfect reference ranges for the majority of them. We can see this is not always the case. We remind that the first two measures accumulate deviations in all “directions”. The KL divergences approach 3 sigma high level for LRG and ELG post-recon, while for BGS they are even further from perfect. are high with a larger scatter for BGS. In the reduced chi-squared, which captures the overall “scaling” with higher accuracy, the mean values for the RascalC runs are shifted significantly for LRG and ELG post-recon, and the mock-to-mock scatter is high in all the cases.
@@
We note that the abovementioned differences are relatively small — in the reduced chi-squared at most for LRG post-recon in the widest range, and in – no more than a percent or two on top of caused by the finite sample size. We should ask whether we trust the realism of the mocks to that level in all aspects of the correlation function multipoles. Moreover, for the real survey matching the clustering between data and simulations will become an additional issue for mocks. We remind that here we are running RascalC with clustering measured directly with each respective mock.
5.4 Model parameter spaces
Now we proceed to project the covariance matrices to model parameters (as described in Section 2.3).
@@ 5.4.1 BAO
In Table 7 we compare the covariances projected to the BAO parameters, the scales and 141414In the Fisher approximation we adopt, the analogs for and should have the same comparison measures.. The comparison measures look consistent with the perfect reference case. The most significant deviations are seen in BGS pre-recon: both the KL divergence and are high, while the reduced chi-squared is almost 3 sigma low on average (meaning that the RascalC covariance is “larger” than the mock sample one). The mock-to-mock scatter in RascalC results is always lower than the noise expected from the finite mock sample size.
We have also plotted the errorbars on and against each other in Fig. 2. They corroborate with Table 7: most deviations are within the 99.7% contour, except a few of the BGS pre-recon runs. However, remember that these plots do not show the covariance of the parameters, which is taken into account with KL divergence and .
5.4.2 Full shape: ShapeFit and direct fit
@@
Additionally, in Fig. 3 we compare the projected errorbars for the scale parameters, and . All the points fall within the 99.7% confidence region in this 2D space, although the other parameters and cross-correlations are ignored in this picture, unlike in Table 8.
@@
Finally, in Table 9 we provide the comparison results for the covariances projected to the direct fit parameters, , , and . We see deviations exceeding 3 sigma in most of the measures for LRG pre-recon and of BGS; LRG post-recon and ELG look consistent.
Still, the differences we see are at a few percent level, and the limitations of the realism of the mocks might be responsible for them.
@@ 6 Conclusions
We have presented and validated the DESI DR1 pipeline for the semi-analytical covariance matrices of the galaxy 2-point correlation functions on realistic mock catalogs.
@@
We have developed a streamlined procedure for the estimation of the semi-analytical covariance matrices for Legendre moments of the 2PCF in separation bins with the RascalC code [28]. The previous implementation [29] required an additional computation with angular bins to mimic the non-Gaussian effects by calibrating the shot-noise rescaling value on the jackknife covariance matrix estimate. Now this calibration can be achieved with one run of the code. This allowed for more efficient massive production of covariance matrices for all the tracers, redshift bins, and galactic caps of DESI DR1 galaxies and quasars data.
@@
We have applied this updated pipeline to an isolated mock catalog with fast (approximate) fiber assignment, representative of DESI DR1, 10 times for 3 selected tracers and redshift bins (LRG3, ELG2 and BGS according to [7]), without and with standard BAO reconstruction applied.
@@
First, we note the difference between the shot-noise rescaling values obtained for the mocks with fast fiber assignment and the data in Section 5.2. However, the parameter calibration on jackknife and mock sample covariance yields very close results in the mock runs (Appendix C). Therefore, this indicates that the fast fiber assignment likely does not reproduce all the properties of the real one relevant to the covariance matrix properties.
@@
Then, we apply the set of compact measures of covariance matrix similarity from [31] in the observable space (correlation function multipoles) in Section 5.3, as well as projected linearly (through Fisher matrix formalism) to the parameters of different models in Section 5.4. We find some deviations that can not be explained solely by the finite sample size limiting the accuracy of the mock-based covariance. However, they are at a few percent level. We argue that our goal is not necessarily to match the mocks perfectly, given their unavoidably approximate nature (including the fast fiber assignment mentioned above). One should also remember that the mocks have a disadvantage when it comes to calibration to the data clustering, especially with blinding.
@@
Since the key application of RascalC covariance matrices at the moment is to the DESI DR1 baryon acoustic oscillations analysis [7], the results projected for the BAO model (Section 5.4.1) are of particular importance. There, for the errorbars of the scale parameters, and , we find a close agreement (, while the standard deviation of the errorbar expected from 1000 mock samples is ), except for BGS before reconstruction.
@@
The covariance matrices for the BGS sample are found to be less consistent in most of the comparisons (except in ShapeFit parameters). We expected them to be more challenging to RascalC because of higher number density and thus higher significance of the 4-point term compared to the 3- and 2-point terms. This already caused slower convergence and could further demonstrate the limitations of the shot-noise rescaling. However, BGS have been challenging for the EZmocks as well, so their sample covariance is likely to be a less robust reference than for LRG and ELG.
@@
The observable-space results (Section 5.3) may leave an impression that RascalC performance worsened with fiber assignment151515Or due to including not only monopole, but also quadrupole and hexadecapole.. However, the previous RascalC validation for early DESI data [31] used an earlier version of EZmocks without fast fiber assignment. Before focusing on the isotropic BAO scales, the covariance matrices there also showed statistically significant variations from the sample covariance, although they were deemed acceptable as comparable to the scatter in semi-analytical results; in this work we use a stricter interpretation, testing whether every RascalC single-mock result is consistent with the perfect-case reference. In future work, we could repeat the tests on the mocks before fiber assignment with all the tracers (and multipoles) used in this paper.
@@
Another direction for further improvement is to study the dependence of shot-noise rescaling on fiber assignment incompleteness by running RascalC on survey sub-regions with a more uniform number of passes. It is also possible that a prescription for higher-point non-Gaussian correlations with a low number of parameters would give a better consistency with the reference than rescaling the covariance matrix terms in which they are nulled. However, for such precision studies, it is instructive to have extremely reliable, realistic and numerous mocks.
Extensions of the approach beyond the standard 2-point function are possible but require extra work. For example, [29] introduced the covariance matrices for isotropic 3-point correlation functions.
@@
In summary, we have validated RascalC semi-analytical covariance matrices for 2PCF as a very viable alternative to the mock-based ones. Despite the increase in the runtime of the RascalC code (from 50-150161616[31] gave the figure based on the number of hyperthreads, not physical cores. to 500-1500 core-hours, or 4-12 node-hours at NERSC Perlmutter supercomputer, Section 5.1), the method is far faster than calibrating, generating, and processing a suite of mocks numerous enough to give an adequate covariance matrix precision. We have shown that the two methods produce similar results given the requirements of the DESI 2024 BAO analysis. The speed advantage of the semi-analytic method permits easier exploration of situations where one cannot afford to regenerate mock catalogs, such as different assumptions about cosmology, galaxy-halo connection, or non-standard sample selections. We therefore expect that such semi-analytic methods can be of broad utility for computing large-scale covariance matrices in wide-field surveys.
@@ Acknowledgements
The data used in this analysis will be made public along with the DESI Data Release 1 (details in https://data.desi.lbl.gov/doc/releases/). The code used in the paper is publicly available at https://github.com/oliverphilcox/RascalC and https://github.com/misharash/RascalC-scripts. The data from the plots, the full set of computed comparison measures and shot-noise rescaling values are available at doi:10.5281/zenodo.10895161 [54].
Appendix A Details on previous covariance estimators
A.1 Full
The full multi-tracer expression for the model covariance (not used in this work, but employed for overlapping tracers’ cross-correlation in [35]) is

### v3

@@ 6.2 Shot-noise rescaling values
| LRG pre-recon | 0.865 | 0.953 | ||
| LRG post-recon | 0.836 | 0.969 | ||
| ELG pre-recon | 0.687 | 0.718 | ||
| ELG post-recon | 0.693 | 0.737 | ||
| BGS pre-recon | 0.855 | 0.897 | ||
| BGS post-recon | 0.872 | 0.934 | ||
Our method uses a rescaling of the shot noise contribution to account for differences between the true small-scale contributions and our Gaussian approximation, as we pointed out in Section 2.1. After reaching a relatively uniform convergence level in the previous section, we should investigate the values of the shot-noise rescaling parameter.
Table 2 shows the shot-noise rescaling values obtained for mocks and DESI DR1 data according to the fiducial procedure (Fig. 1). Interestingly, we find that all the shot-noise rescaling values are smaller than 1; i.e., the jackknife variations are smaller than predicted by Gaussian approximation with standard Poisson shot noise. Previous mock studies [28, 29, 30, 35], on the contrary, obtained shot-noise rescaling values greater than 1.
We also see that the shot-noise rescaling values are significantly lower for mocks than for the data. The difference is most pronounced for ELGs, which have the lowest shot-noise rescaling values for both mocks and data. ELGs are also impacted most by the fiber incompleteness effect due to their lower priority compared to other dark-time targets, LRG and QSO [70]. This suggests that the low shot-noise rescaling is related to fiber assignment.
To test whether fiber assignment is the main cause for low and different shot-noise rescaling values, we have performed additional RascalC computations using more realistic Abacus mocks and the DESI fiber assignment algorithm (AMTL). We only used one realization for LRG and ELG in each case to save computing time.
| Data | AMTL (correct) | FFA (approximate) | Complete | |||
| Abacus | EZmocks | Abacus | EZmocks | Abacus | ||
| LRG NGC | 0.865 | 0.845 | 0.886 | 0.721 | 0.758 | 0.934 |
| LRG SGC | 0.953 | 0.954 | 0.961 | 0.781 | 0.784 | 0.962 |
| ELG NGC | 0.687 | 0.649 | 0.672 | 0.378 | 0.380 | 0.965 |
| ELG SGC | 0.718 | 0.707 | 0.744 | 0.405 | 0.403 | 0.968 |
Table 3 shows the shot-noise rescaling values for data and different mocks (Abacus or EZmocks) with different fiber assignment models (FFA or AMTL). We also include an Abacus mock before fiber assignment (“complete”). We see that the fiber assignment modeling method makes a bigger difference than the type of the mocks. The approximate fast fiber assignment gives the lowest shot-noise rescaling values. The DESI fiber assignment algorithm applied to the mocks (AMTL) gives very similar to the data. The complete Abacus mocks have larger shot-noise rescaling values than data and fiber-assigned mocks. This confirms the fiber assignment as a key factor affecting the parameter.
The discrepancy in shot-noise rescaling values also raises concerns about the quality of approximations in the FFA algorithm. However, of all the mock types only EZmocks processed with FFA are numerous enough for a precise sample covariance estimate. Therefore we continue using them in the remainder of this paper.
| NGC | SGC | |||
| Jackknife | Mock sample | Jackknife | Mock sample | |
| LRG pre | ||||
| LRG post | ||||
| ELG pre | ||||
| ELG post | ||||
| BGS pre | ||||
| BGS post | ||||
@@
We performed the final consistency tests of the shot-noise optimization procedure in light of the concerns about fiber assignment and jackknife discussed in Section 4.2. We optimized the shot-noise rescaling based on mock sample covariance (the process is shown schematically in Fig. 2). We show the resulting values along with the baseline, jackknife-based ones in Table 4, and find them very close for all cases. In other words, calibration of shot-noise rescaling on jackknives still brings us close to an optimal fit on the mocks. With that, we have decided to proceed further with the validation process using the jackknife-calibrated shot-noise rescaling values. This will show how close this nearly optimal fit is to the mock sample covariance.
7 Comparison between semi-analytical and mock sample covariance matrices
Thus we reach the last validation part: comparison of the semi-analytical covariances obtained from single mock catalogs with the mock sample covariance matrices. We compute three similarity measures (Section 2.2) between these covariance matrices for each tracer before and after standard BAO reconstruction (pre- and post-recon). Instead of presenting 10 numbers (corresponding to each RascalC realization) for each case in the paper, we provide their mean and standard deviation. The full set of comparison measure values is available in the supplementary material292929doi:10.5281/zenodo.10895161. We also provide the “perfect” reference row, listing the statistical properties of the similarity measures between the true covariance and a mock sample covariance matrix based on their dimension (also obtained in Section 2.2).
We note that the standard deviation in the “perfect” row is different and independent from the others. The “perfect” standard deviation characterizes the distribution of the random difference between the sample covariance estimate from 1000 realizations and the true covariance. In each other row (tracer + pre- or post-recon combination) the sample covariance estimate is fixed, and the standard deviation describes the scatter resulting from 10 different RascalC covariance realizations.
Ideally, every RascalC covariance matrix would be consistent with the true covariance. To see whether this is the case, we compute the difference between each comparison measure for each RascalC covariance realization and the corresponding “perfect” mean in “perfect” standard deviations. We summarize the 10 resulting quantities by the mean and standard deviation as well and provide them below the summary of the similarity measure itself. This sets the common structure for all tables in this section (i.e., Tables 5, 6, 7, 8, 9 and 10.).
7.1 Observables
We begin with the covariances for correlation function multipoles, in other words, in the space of observables.
@@ 7.1 Observables
Table 5 shows the comparison measures for the full covariance matrices, covering radial bins for all three multipoles. We can see that some RascalC results deviate significantly from the “perfect” (i.e., the true covariance behavior). We remind that the KL divergence and accumulate deviations in all “directions”. The KL divergences exceed the ideal expectation value by nearly for LRG and ELG post-recon, whereas for BGS they are even further from perfect. are high with a larger scatter for BGS. In the reduced chi-squared, which captures the overall “scaling” with higher accuracy, the mean values for the RascalC runs are shifted significantly for LRG and ELG post-recon, and the mock-to-mock scatter is high in all the cases.
@@
We note that the abovementioned differences are relatively small. In the reduced chi-squared, they are at most for LRG post-recon in the widest range, and in – no more than a percent or two on top of caused by the finite sample size. We should ask whether we trust the realism of the mocks to such a high level in all aspects of the correlation function multipoles. Moreover, for the real survey matching the clustering between data and simulations will become an additional issue for mocks. On the other hand, our RascalC computations use the correlation function measured directly from single mock catalogs similar to real data. Thus we conclude that the semi-analytic method performance is very compelling.
7.2 Model parameters
We proceed to project the covariance matrices (as described in Section 2.3) to the parameters of the models listed in Section 4.4.
We use the same model Jacobian (Eq. 2.20) for projecting RascalC covariances (Eq. 2.19) and the mock sample covariances (Eq. 2.21) into the parameter space. We compute the partial derivatives at the best fit of each model to the mean clustering of all the available mocks for each galaxy type. Since BAO reconstruction changes clustering, we compute separate Jacobians before and after reconstruction. We use the mock sample covariance matrix in this fit. Using different Jacobian matrices for each covariance matrix would complicate the comparison.
A more thorough investigation of the covariance matrix impact on the model fits is presented in the companion paper [36]. They perform fits to each mock clustering using different covariance matrices (mock sample and semi-analytical) and compare the resulting best parameter values as well as errorbar estimates. On the flip side, with such a detailed approach [36] can not test multiple RascalC single-mock covariances.
@@ 7.2.1 BAO
In Table 8 we compare the covariances projected to the BAO parameters, the scaling parameters and 303030Because we use Fisher matrix formalism, the covariance matrices projected for and should have the same comparison measures.. The comparison measures look consistent with the perfect reference case. The most significant deviations are seen in BGS pre-recon: both the KL divergence and are high, while the reduced chi-squared is almost 3 sigma low on average (meaning that the RascalC covariance is “larger” than the mock sample one). The mock-to-mock scatter in RascalC results is always lower than the noise expected from the finite mock sample size.
We have also plotted the errorbars on and against each other in Fig. 3. They corroborate Table 8. Almost all deviations are within the 99.7% contour. The only exceptions are two BGS pre-recon realizations. However, we note that these plots do not show the covariance of the parameters, which is taken into account with KL divergence and .
7.2.2 Full shape: ShapeFit and full modeling
@@
Additionally, in Fig. 4 we compare the projected errorbars for the scaling parameters, and . All the points fall within the 99.7% confidence region in this 2D space, although the other parameters and cross-correlations are ignored in this picture, unlike in Table 9.
@@
Finally, in Table 10 we provide the comparison results for the covariances projected to the full modeling parameters, , , and . We see deviations exceeding 3 sigma in most of the measures for LRG pre-recon and of BGS; LRG post-recon and ELG look consistent. Still, the differences we see are at a few percent level, and may be caused by limited realism of the mocks.
@@ 8 Conclusions
We present and validate the DESI DR1 pipeline for the semi-analytical covariance matrices of the galaxy 2-point correlation functions on realistic mock catalogs including a model of fiber assignment.
@@
We develop a streamlined procedure for the estimation of the semi-analytical covariance matrices for Legendre moments of the 2PCF in separation bins with the RascalC code [30]. The previous implementation [31] required an additional computation with angular bins to mimic the non-Gaussian effects by calibrating the shot-noise rescaling value on the jackknife covariance matrix estimate. Now we can perform this calibration together with the construction of the covariance model for Legendre moments. This allowed for more efficient massive production of covariance matrices for all the tracers, redshift bins, and galactic caps of DESI DR1 galaxies and quasars data.
@@
We apply this updated pipeline to mock catalogs with fast (approximate) fiber assignment, representative of DESI DR1. We cover 3 selected tracers and redshift bins (LRG3, ELG2 and BGS according to [7]), without and with standard BAO reconstruction applied. We use a single mock catalog and repeat the procedure for 10 different realizations to assess the impact of realistic fluctuations in the input quantities.
@@
First, we note the difference between the shot-noise rescaling values obtained for the mocks with fast fiber assignment and the data in Section 6.2. We show that the fiber assignment modeling (and the associated weighting scheme) is the main factor causing the difference. Then we find the parameter calibration on jackknife and mock sample covariance yields very close results in the mock runs. This motivates us to proceed with the comparison of the covariance matrices.
@@
Then, we apply the set of compact measures of covariance matrix similarity from [35]. We use covariances in the observable space (correlation function multipoles) in Section 7.1, as well as projected linearly (through Fisher matrix formalism) to the parameters of different models in Section 7.2. We find some deviations that cannot be explained solely by the finite sample size limiting the accuracy of the mock-based covariance. However, these differences are at a few percent level.
We argue that this level of agreement is sufficient for real-world applications. First, mocks necessarily involve approximations to make sufficiently many catalogs. Second, simulations are never perfectly representative of data in terms of clustering due to the limited precision of the measurements. Blinding can aggravate this issue by forcing the mock tuning to rely on earlier, smaller samples with larger uncertainties in their clustering. Additionally, some simulations, including EZmocks [7, 15], are only matched to the 1- and 2-point statistics of the data, whereas the covariance matrices are impacted by 3- and 4-point functions. Therefore matching the mocks perfectly is not necessarily a reasonable goal.
@@
Our results for the BAO model (Section 7.2.1) are particularly important because RascalC covariance matrices are used in the DESI DR1 baryon acoustic oscillations analysis [7]. The errorbars of the scale parameters ( and ) predicted from RascalC agree to with the mock sample covariance. The standard deviation expected from the sample covariance of 1000 mocks itself is . When we exclude BGS before reconstruction (not used for the main DESI DR1 BAO), the errorbars agree within . Therefore we report a close match between the semi-analytic and mock sample covariance.
@@
We find the covariance matrices for the BGS sample less consistent in most comparisons313131A notable exception is ShapeFit parameters in Section 7.2.2.. We expected them to be more challenging for RascalC because of higher number density and thus higher significance of the 4-point term compared to the 3- and 2-point terms. This already caused slower convergence and could further demonstrate the limitations of the shot-noise rescaling. However, BGS have been challenging for the EZmocks as well [15]323232Moreover, DESI BGS EZmocks had not been used in [34]., so their sample covariance is likely to be a less robust reference than for LRG and ELG.
@@
The observable-space results (Section 7.1) may leave an impression that RascalC performance worsened with fiber assignment333333Or due to including not only monopole, but also quadrupole and hexadecapole.. However, we think the change of interpretation primarily causes this. The previous RascalC validation for early DESI data [35] used an earlier version of EZmocks without any fiber assignment effects. The covariance matrices there also showed statistically significant variations from the sample covariance in observable space. They were deemed acceptable as comparable to the scatter in semi-analytical results. In this work we use a stricter interpretation, testing whether every RascalC single-mock result is consistent with the perfect-case reference. In future work, we could repeat the tests on the mocks before fiber assignment with all the tracers (and multipoles) used in this paper.
@@
Another direction for further improvement is to study the dependence of shot-noise rescaling on fiber assignment incompleteness. This could be achieved by running RascalC on survey sub-regions with different completeness patterns, set by the number of passes. It is also possible that a prescription for higher-point non-Gaussian correlations with a low number of parameters would give a better consistency with the reference than rescaling the covariance matrix terms in which they are nulled. However, for such precision studies, it is instructive to have extremely reliable, realistic and numerous mocks.
The semi-analytic approach can extend beyond the standard 2-point function. First, the cross-covariances of different tracers are provided in Appendix A. The full cross-covariance has more bins and requires more mocks than a single covariance for validation at similar precision. Consistent simulated catalogs for different tracers are also crucial for capturing realistic cross-correlations. Second, [72] have introduced the covariance matrices for the modified power spectra, including the cross-covariance with correlation functions. Third, [31] have derived the covariance matrices for isotropic 3-point correlation functions. It involves higher-order correlation functions up to 6 points. Approximating all of them in fast mocks is challenging. Thus these extensions require extra work, particularly on the mock side for validation.
@@
In summary, we have confirmed RascalC semi-analytical covariance matrices for 2PCF as a very viable alternative to the mock-based ones. Despite the increase in the runtime of the RascalC code (from 100-300 [35] to 500-1500343434See Section 6.1. Potentially times smaller using the later profiling results. core-hours), the method is far faster than calibrating, generating, and processing a suite of mocks numerous enough to give an adequate covariance matrix precision. We have shown that the two methods produce similar results given the requirements of the DESI 2024 BAO analysis. The speed advantage of the semi-analytic method permits easier exploration of situations where one cannot afford to regenerate mock catalogs, such as different assumptions about cosmology, galaxy-halo connection, or non-standard sample selections. We therefore expect that such semi-analytic methods can be of broad utility for computing large-scale covariance matrices in wide-field surveys.
@@ Acknowledgements
This work has used the following software packages: astropy [73, 74, 75], Jupyter [76, 77], matplotlib [78], numpy [79], pycorr [33, 80, 81], python [82], scipy [83, 84], and scikit-learn [85, 86, 87].
This research has made use of NASA’s Astrophysics Data System. Software citation information has been aggregated using The Software Citation Station [88, 89].
@@
The data used in this analysis will be made public along with the DESI Data Release 1 (details in https://data.desi.lbl.gov/doc/releases/). The code used in the paper is publicly available at https://github.com/oliverphilcox/RascalC and https://github.com/cosmodesi/RascalC-scripts. The data from the plots, the full set of computed comparison measures and shot-noise rescaling values are available at doi:10.5281/zenodo.10895161 [90].
Appendix A Estimators for multi-tracer covariance matrices
In this section, we provide the generalized multi-tracer expressions for reference. We have not endeavored to validate them in this work because in the EZmocks suite different tracers have mostly been made from boxes at different redshifts. We expect that the shift of the galaxies’ positions with time will not allow self-consistent cross-correlations in such circumstances. DESI has primarily focused on the LRG and ELG overlap in bin [39], but EZmocks for these tracers have been made from box snapshots at redshifts 1.1 and 0.95 respectively.
We add one capital Latin superscript to all the overdensities to signify the tracer involved (e.g. ). For counts, it is two letters accordingly (e.g. ).
A.1 Angular bins
The Landy-Szalay estimator for the multi-tracer correlation function is
We make the following shorthand for the multi-tracer covariance:
It is important to note that the shot-noise approximation works only for the same-tracer overdensities:
The full multi-tracer expression for the model covariance is from [30] but with some changes in notation: